Let P(x) be a polynomial of degree 5 such that P(k)=3^k for k = 0,1,2,3,4,5. Then the sum of digit in P(6) is equal to???

Grade:12

1 Answers

Arun
6 years ago
As we know
P(k) = 3^k
Hence
P(6) = 3^6
P(6) = 729
Thus sum of digit in 729 = 7+2+9 = 18
 
Hope it helps